OpenStudy (anonymous):

Solve 5y + 3 = 2(y – 1). y = -5/3 y = -5/7 y = -4/3 y = -4/7

OpenStudy (radar):

5y+3=2y-2 5y-2y=-2-3 3y=-5 y=-5/3

OpenStudy (anonymous):

yea radar is correct that is the answer

OpenStudy (anonymous):

aren't you glad you know?